# TrafficGamer: Reliable and Flexible Traffic Simulation for Safety-Critical Scenarios with Game-Theoretic Oracles

### v1

@@ Results
To better evaluate the effectiveness of our method in replicating the driving styles of human drivers across various scenarios, we categorize the scenarios into specific types, including fork road and cross road. Figure 4 (rows from 3 to 6) compares the learned and actual distributions of distance and speed. In the crossroad scenario (rows 5 and 6 in Figure 4), TrafficGamer  achieves performance comparable to, albeit slightly less robust than QCNet in maintaining fidelity. This result is anticipated since QCNet is designed to closely mimic the observed behaviors of drivers, resulting in trajectories that heavily overlap with those in the source dataset. Another key observation is that our method, TrafficGamer, significantly outperforms standard MARL approaches such as MAPPO. This demonstrates TrafficGamer’s capability to accurately replicate a wide range of human driving behaviors and generate realistic driving scenarios. We also check the fidelity in the Waymo motion dataset, and the results closely match the ground-truth distribution. More details can be found in supplementary Section 3d.
@@
Table 1: The mean $\pm$ std of the CCE-gap, calculated using the first method, is based on an average of over 300 episodes. The best average performance is highlighted in bold. N/A indicates this agent does not exist in this scenario.
@@
QCNet 5.59$\pm$1.35 16.46$\pm$3.66 11.50$\pm$0.38 6.43$\pm$0.66 18.16$\pm$2.57 20.58$\pm$0.48 13.53$\pm$0.30
GameFormer 15.88$\pm$4.28 51.62$\pm$14.47 20.96$\pm$3.78 29.89$\pm$9.07 13.76$\pm$1.73 25.27$\pm$0.79 14.93$\pm$1.11
MAPPO 6.21$\pm$0.60 8.60$\pm$0.50 9.85$\pm$0.78 9.69$\pm$1.01 18.37$\pm$0.91 8.46$\pm$1.08 6.01$\pm$1.14
TrafficGamer 2.10$\pm$0.40 2.71$\pm$0.79 3.03$\pm$0.31 2.27$\pm$0.69 1.97$\pm$0.30 1.30$\pm$0.40 2.14$\pm$0.28
@@
Table 2: The mean $\pm$ std of the CCE-gap, calculated using the second method, is based on an average of over 300 episodes. The best average performance is highlighted in bold. N/A indicates this agent does not exist in this scenario.
@@
Unlike standard RL algorithms that focus on reward maximization, in the multi-agent autonomous driving environment, we primarily consider exploitability [43], which measures the extent to which a vehicle’s policy can exploit the current traffic policies. An ideal driving equilibrium should have zero exploitability, meaning no single vehicle can achieve greater benefits by continually improving the policy. In this work, we follow the methodology outlined in [47] and utilize the CCE-gap (see definition 1) to measure exploitability. Unlike the common-studied toy environment (e.g., Bargaining, TradeComm, and Battleship [44]), the traffic scenarios involve complex game contexts, multiple agents, and continuous action space, which makes the best response $\pi^{\dagger}_{i}$ computationally intractable, and accurately calculating the exact CCE-gap becomes challenging. Therefore, we estimate the CCE-gap by empirical approximating $\pi^{\dagger}_{i}$ with the following methods:
@@
  • Breaking the Equilibrium. Upon the studied algorithm converges, we estimate each agent’s best response under the current equilibrium, denoted as $\pi^{\dagger}_{i}$, by fixing the policies of the other $I-1$ agents and continuously encouraging this agent to maximize its current reward. This process assesses the agent’s ability to disrupt the existing equilibrium. If none of the policies can achieve significantly higher rewards, it indicates that the experimental algorithm has successfully identified a reliable CCE.
@@
  • Restricting Action Space. To overcome the computational intractability caused by complex driving behaviors, we constrain the decision-making domain to choose actions from a predefined candidate action set. This set is generated by a pre-trained action predictor (see objective 9), which identifies and ranks the top ${K}$ actions most similar to those observed in the dataset. Within this tractable action space, we compute $\pi^{\dagger}_{i}$ by selecting the actions that optimally maximize rewards at each step.
@@
We compare the performance of TrafficGamer in all Argoverse 2 scenarios with other baselines based on the aforementioned CCE-Gap. Figure 5 and Figure 6 present the CCE-gap of each agent during training. The results show that the training curves of QCNet and GameFormer perform unstable and struggle to converge to the CCE. It illustrates that end-to-end imitation learning methods can not model competitiveness between agents, which makes it difficult to optimize policies for capturing CCE. MAPPO performs better than other baselines, but it still falls short of the results obtained by TrafficGamer. MAPPO faces challenges in efficiently exploring the entire policy space of the Multi-agent game environment during the optimization process, making it difficult to capture the underlying CCE. As a solver defined for CCE, TrafficGamer ensures that the agents’ policies are distributionally aligned with human-driven policies and supports stable exploration. This allows each agent to learn the optimal policy and gradually converge to an approximate CCE.
@@
Table 1 and Table 2 present the CCE-gap of each agent at the end of training. TrafficGamer achieves a smaller CCE-gap compared to other methods, demonstrating its superior exploitability across various scenarios. While TrafficGamer achieves a smaller CCE-gap compared to other methods, it exhibits slightly reduced performance for certain agents in some scenarios. An in-depth analysis reveals two main reasons for this phenomenon: 1) TrafficGamer may struggle to effectively adapt to rapid changes in opponents’ strategies under specific conditions. For example, when surrounding vehicles accelerate, a particular vehicle may need to decelerate to maintain safe spacing. 2) The presence of other vehicles significantly influences the decision-making processes of algorithm-controlled vehicles. For example, if a car on the other hand begins to merge into the current lane, the algorithm might adopt a more conservative driving strategy to prioritize safety, thereby influencing exploiting. Overall, in a multi-agent game environment, it is essential to evaluate the collective performance of all agents. TrafficGamer is capable of learning a CCE that approximates the optimal solution, resulting in a more comprehensive modeling of vehicle driving behaviors.
@@
Figure 5: CCE-gap obtained from the first calculation approach, where each row represents a scenario and scenarios 1-6 correspond to Y-Junction, Dense-lane intersection, Roundabout, Dual-lane intersection, T-Junction, and Merge respectively. Each column corresponds to one of the agents in the multi-agent environment. Because of the space limit, the CCE-gap results of other agents can be seen in Supplementary Section 3a.
@@
Figure 6: CCE-gap obtained from the second calculation approach, where each row represents a scenario and scenarios 1-6 correspond to Y-Junction, Dense-lane intersection, Roundabout, Dual-lane intersection, T-Junction, and Merge respectively. Each column corresponds to one of the agents in the multi-agent environment. Because of the space limit, the CCE-gap results of other agents can be seen in Supplementary Section 3a.
Figure 7: Variation of Lagrangian Parameters under varying constraints, where each row represents a scenario and scenarios 1-6 correspond to Y-Junction, Dense-lane intersection, Roundabout, Dual-lane intersection, T-Junction, and Merge respectively. Each column corresponds to one of the agents in the multi-agent environment. Because of the space limit, the lagrangian parameter results of other agents can be seen in Supplementary Section 3b.
@@
To demonstrate the ability of TrafficGamer to generate diverse traffic scenarios, we visualize the generated scenarios featuring varying degrees of competitive behavior. However, automated driving scenarios involve complex elements including road structures, traffic regulations, and vehicle behaviors, which can be challenging to interpret directly from raw data. To address this, we simulate the actual behavior of vehicles in various safety-critical scenarios and traffic conditions using 2D and 3D visualizations. These are captured in third-person and first-person views, to provide a clearer understanding of the dynamics at play. To facilitate 2D visualization, TrafficGamer supports the display of lines, markers, the road network, vehicle positions, and movement trajectories on a unified 2D plane by following the formats established by Argoverse 2 [45] and Waymo [46]. This approach uses concise symbols to represent various traffic elements, enhancing the intuitiveness of scene analysis. To capture the real-world complexity of traffic behaviors more accurately, TrafficGamer enables large-scale 3D traffic scenario modeling and simulation through ScenarioNet [54], based on the MetaDrive simulator [7]. The scenarios generated can be replayed and explored interactively, ranging from Bird’s Eye View layouts to realistic 3D renderings in ScenarioNet.
@@
In specific, this experiment explores how well TrafficGamer can facilitate the generation of different safety-critical scenarios under twelve distinct traffic scenarios. To generate diverse traffic scenarios, we design the constrained and risk-sensitive policy optimization [55, 56] for capturing the equilibrium subject to different levels of tightness. Specifically, by adjusting the inter-vehicle distance constraints and risk coefficients, we derive a diverse number of traffic scenarios under different game contexts. Figure 8 illustrates the details of our results.
@@
By comparing the scenarios generated with different levels of inter-vehicle distance constraints (from top to bottom in Figure 8 and Figure 9) , we find as the inter-vehicle distance increases, our model adapts, leading to traffic scenarios characterized by less competitive behavior and safer navigation across all examined situations. This outcome arises because maintaining distance constraints requires the cooperation of multiple vehicles. Imposing more restrictive constraints (i.e., increasing the distance) significantly enhances the impact of this cooperation on the optimization objective 16 (dynamically controlled by the Lagrange parameter $\lambda$). Figure 7 displays the variation of the Lagrangian penalty factor during the training process. As the constraints become tighter, the Lagrangian penalty term also increases, indicating that it has a larger impact on the objective 16. On the other hand, if we reduce the required distance, agents begin to prioritize their interests, which significantly increases the likelihood of traffic congestion where no agent can further optimize their policy. All these dynamic scenarios are characterized by the learned CCEs.
@@
Similarly, by comparing the scenarios generated with different levels of risk sensitivity (from left to right in Figure 8), we find imposing a higher confidence level leads to more conservative and cautious driving behaviors. A higher confidence level forces the agent to satisfy the constraint with greater probability, resulting in driving strategies that feature lower speeds, increased spacing between vehicles, and more careful navigation through complex traffic scenarios. On the other hand, setting a lower confidence level results in more aggressive and risk-seeking driving behaviors, characterized by faster vehicle speeds and shorter following distances. The system’s tolerance for some aggressive behaviors (such as overtaking etc.) has increased. This approach can lead to a higher risk of collisions and increase the likelihood of critical safety scenarios in traffic.
@@
These diverse scenarios are crucial for investigating the trade-offs between aggressive and conservative driving behaviors. By analyzing AV policies within these scenarios, we can more effectively evaluate the optimality of driving strategies across different levels of competition and congestion. Furthermore, the comprehensive 2D and 3D simulations provided by our TrafficGamer  from both third-person and first-person perspectives offer a detailed understanding of traffic dynamics. These simulations demonstrate how variations in risk sensitivity and distance constraints impact vehicle behavior in real-world driving scenarios.
@@
Figure 8: Visualization of generated trajectories with high diversity (Argoverse2). We visualize all Argoverse2 scenarios in a 3x3 grid layout with 2D and 3D simulations. In each grid, the distance constraint relaxes from top to bottom, and the risk level strengthens from left to right. In the 2D scenes, yellow cars are algorithm-controlled, green cars and gray cars are environmental vehicles, black lines are vehicle trajectories, and blue arrows indicate travel direction. In the 3D scenes, red cars are algorithm-controlled, while blue cars represent environmental vehicles.
@@
Figure 9: Visualization of generated trajectories with high diversity (Waymo). We visualize all Waymo scenarios in a 3x3 grid layout with 2D and 3D simulations. In each grid, the distance constraint relaxes from top to bottom, and the risk level strengthens from left to right. In the 2D scenes, yellow cars are algorithm-controlled, green cars and gray cars are environmental vehicles, black lines are vehicle trajectories, and blue arrows indicate travel direction. In the 3D scenes, red cars are algorithm-controlled, while other cars represent environmental vehicles.
@@ Discussion
We demonstrate that our model, TrafficGamer, can effectively represent various safe-critical traffic scenarios by capturing a range of CCEs. To the best of our knowledge, this is the first algorithm that fine-tunes generative world models to accurately model competitive and collaborative behaviors among multiple agents across varying degrees of constraints and dynamic environments. Most importantly, TrafficGamer can accurately characterize traffic congestion scenarios frequently observed in reality but are underrepresented in datasets, such as roundabouts, intersections, and merge points. This capability ensures the high fidelity of TrafficGamer in real-world applications.
To facilitate the reliable generation of our safety-critical scenarios, we resolve three critical challenges, including 1) how to guarantee the fidelity of generated trajectories, 2) how to efficiently capture the CCE of each scenario by modeling the competition behaviors of vehicles, and 3) how to dynamically adapt the strength of the equilibrium. These safety-critical scenarios can serve as important testbeds for AV controlling algorithms evaluating the reliability and robustness of AV control algorithms before their practical deployment. Traditional simulation systems have relied heavily on manually designed rules or data-driven trajectory imitation, often resulting in scenarios that lack fidelity and diversity. Our model, TrafficGamer, addresses these limitations by generating a variety of realistic, safety-critical scenarios. It is important to note that our method currently models vehicle behaviors on a static map. Future enhancements for TrafficGamer may include integrating multi-modal large models to generate scenarios that account for varying vehicle behaviors in response to environmental factors, such as weather conditions, time of day, and topography.

### v2

@@ Results
To better evaluate our method’s effectiveness in replicating humans’ driving styles across various scenarios, we categorize the scenarios into specific types, including fork road and cross road. We then compare the learned and actual distributions of distance and speed (Figure 4, rows from 3 to 6). In the crossroad scenario (rows 5 and 6 in Figure 4), TrafficGamer ’s performance is comparable to, albeit slightly less robust than, that of QCNet in maintaining fidelity. This result is as anticipated, as QCNet is designed to closely mimic the observed behaviors of drivers, resulting in trajectories that heavily overlap with those in the source dataset. Another key observation is that our method, TrafficGamer, significantly outperforms standard MARL approaches such as MAPPO. This demonstrates TrafficGamer’s capability in accurately replicating a wide range of human driving behaviors and generating realistic driving scenarios. We also checked the fidelity of TrafficGamer for the Waymo motion dataset, and the results closely match the ground-truth distribution. More details can be found in Supplementary Section 3d.
@@
Table 1: CCE-gap value at the last time step calculated using the Breaking the Equilibrium method
@@
QCNet 5.59$\pm$1.35 16.46$\pm$3.66 11.50$\pm$0.38 6.43$\pm$0.66 17.16$\pm$2.57 20.58$\pm$0.48 13.53$\pm$0.30
GameFormer 15.88$\pm$4.28 51.62$\pm$14.47 20.96$\pm$3.78 29.89$\pm$9.07 12.76$\pm$1.73 25.27$\pm$0.79 14.93$\pm$1.11
MAPPO 2.21$\pm$0.60 4.60$\pm$0.50 5.85$\pm$0.78 5.69$\pm$1.01 13.37$\pm$0.91 4.46$\pm$1.08 6.01$\pm$1.14
TrafficGamer 2.10$\pm$0.40 2.71$\pm$0.79 3.03$\pm$0.31 2.27$\pm$0.69 0.97$\pm$0.30 1.30$\pm$0.40 2.14$\pm$0.28
@@
Table 2: CCE-gap value at the last time step calculated using the Restricting Action Space method
@@
Standard RL algorithms focus on reward maximization; however, in the multi-agent autonomous driving environment, we primarily consider exploitability [43], which measures the extent to which a vehicle’s policy can exploit the current traffic policies. An ideal driving equilibrium should have zero exploitability, meaning no single vehicle can achieve greater benefits by continually improving the policy. In this paper, we follow the methodology outlined in [47] and utilize the CCE-gap (see Definition 1) to measure exploitability. Unlike the commonly studied toy environments (e.g., Bargaining, TradeComm, and Battleship [44]), the traffic scenarios involve complex game contexts, multiple agents, and continuous action space, which makes the best response $\pi^{\dagger}_{i}$ computationally intractable, thus accurately calculating the exact CCE-gap becomes challenging. Therefore, we estimate the CCE-gap by empirically approximating $\pi^{\dagger}_{i}$ via the following methods:
@@
  • Breaking the Equilibrium. Upon the convergence of the studied algorithm converges, we estimate each agent’s best response under the current equilibrium, denoted as $\pi^{\dagger}_{i}$, by fixing the policies of the other $I-1$ agents and continuously encouraging this agent to maximize its current reward. This process assesses the agent’s ability to disrupt the existing equilibrium. If none of the policies can achieve significantly higher rewards, it indicates that the experimental algorithm has successfully identified a reliable CCE.
@@
  • Restricting Action Space. To overcome the computational intractability caused by complex driving behaviors, we constrain the choice of the decision-making domain to choose actions from a predefined candidate set. This set is generated via a pre-trained action predictor (see Objective 9), which identifies and ranks the top ${K}$ actions most similar to those observed in the dataset. Within this tractable action space, we compute $\pi^{\dagger}_{i}$ by selecting the actions that optimally maximize rewards at each step.
@@
We compare the performance of TrafficGamer in all Argoverse 2 scenarios with other baselines based on the aforementioned CCE-Gaps. The CCE-gaps achieved for each agent during training are presented in Figure 5 and Figure 6. The training curves of QCNet and GameFormer show their performance is unstable and they struggle to converge at the CCE. The results illustrates that end-to-end imitation learning methods can not model competitiveness between agents, which makes it difficult to optimize policies for capturing the CCE. MAPPO performs better than other baselines, but it falls short of the results obtained by TrafficGamer. MAPPO faces challenges in efficiently exploring the entire policy space of the multi-agent game environment during the optimization process, making it difficult to capture the underlying CCE. As a solver defined for CCE, TrafficGamer ensures that the agents’ policies are distributionally aligned with human-driven policies and supports stable exploration. This allows each agent to learn the optimal policy and gradually converge to an approximate CCE.
@@
The CCE-gap of each agent at the end of the training is presented in Table 1 and Table 2  [^1] We calculate the mean ± std of the CCE-gap, based on an average of over 300 episodes. The best average performance is highlighted in bold. N/A indicates this agent does not exist in this scenario.. TrafficGamer achieves a smaller CCE-gap than other methods, demonstrating its superior exploitability across various scenarios. However, it exhibits slightly lower performance for certain agents in some scenarios. An in-depth analysis reveals the two main reasons for this phenomenon: (1) TrafficGamer may struggle to effectively adapt to rapid changes in opponents’ strategies under specific conditions. For example, when surrounding vehicles accelerate, a particular vehicle may need to decelerate to maintain safe spacing. (2) Other vehicles’ presence significantly influences algorithm-controlled vehicles’ decision-making processes. For example, if a car on the other hand begins to merge into the current lane, the algorithm might adopt a more conservative driving strategy to prioritize safety, thereby influencing exploiting. Overall, in a multi-agent game environment, it is essential to evaluate the collective performance of all agents. TrafficGamer is capable of learning a CCE that approximates the optimal solution, resulting in more comprehensive modeling of vehicle-driving behaviors.
@@
Figure 5: CCE-gap obtained from the Breaking the Equilibrium approach. Each row represents a scenario and scenarios 1-6 correspond to Y-Junction, Dense-lane intersection, Roundabout, Dual-lane intersection, T-Junction, and Merge respectively. Each column corresponds to one of the agents in the multi-agent environment. Because of the space limitations, the CCE-gap results of other agents can be seen in Supplementary Section 3a.
@@
Figure 6: CCE-gap obtained from the Restricting Action Space approach. Each row represents a scenario and scenarios 1-6 correspond to Y-Junction, Dense-lane intersection, Roundabout, Dual-lane intersection, T-Junction, and Merge respectively. Each column corresponds to one of the agents in the multi-agent environment. Because of the space limitations, the CCE-gap results of other agents can be seen in Supplementary Section 3a.
@@
To demonstrate the ability of TrafficGamer to generate diverse traffic scenarios, we can visualize the generated scenarios with varying degrees of competitive behavior. Automated driving scenarios involve complex elements including road structures, traffic regulations, and vehicle behaviors, which can be challenging to interpret directly from raw data. To address this, we simulate the actual behavior of vehicles in various safety-critical scenarios and traffic conditions using 2D and 3D visualizations. These are captured in third-person and first-person views, to provide a clearer understanding of the dynamics at play. To facilitate 2D visualization, TrafficGamer supports the display of lines, markers, the road network, vehicle positions, and movement trajectories on a unified 2D plane by following the formats established for Argoverse 2 [45] and Waymo [46]. This approach uses concise symbols to represent various traffic elements, enhancing the intuitiveness of scene analysis. To capture the real-world complexity of traffic behaviors more accurately, TrafficGamer enables large-scale 3D traffic scenario modeling and simulation through ScenarioNet [54], based on the MetaDrive simulator [7]. The scenarios generated can be replayed and interactively explored from multiple perspectives, ranging from Bird’s Eye View layouts to realistic 3D renderings in ScenarioNet.
@@
In this experiment, we explore how well TrafficGamer can generate different safety-critical scenarios under 12 distinct traffic scenarios. To generate diverse traffic scenarios, we design a constrained and risk-sensitive policy optimization method [55, 56] for capturing the equilibrium subject to different levels of tightness. Specifically, by adjusting the inter-vehicle distance constraints and risk coefficients, we derive a diverse number of traffic scenarios under different game contexts. Figure 7 illustrates the details of our results.
@@
By comparing the scenarios generated with different levels of inter-vehicle distance constraints (from top to bottom in Figure 7 and Figure 8) , we find that, as the inter-vehicle distance increases, our model adapts, leading to traffic scenarios characterized by less competitive behavior and safer navigation across all examined situations. This outcome arises because maintaining distance constraints requires the cooperation of multiple vehicles. Imposing more restrictive constraints (i.e., increasing the distance) significantly enhances the impact of this cooperation on the optimization of Objective 16 (dynamically controlled by the Lagrange parameter $\lambda$). Figure 9 displays the variation in the Lagrangian penalty factor during the training process. As the constraints become tighter, the Lagrangian penalty term also increases, indicating that it has a larger impact on Objective 16. On the other hand, when we reduce the required distance, agents begin to prioritize their interests, which significantly increases the likelihood of traffic congestion where no agent can further optimize their policy. All these dynamic scenarios are characterized by the learned CCEs.
@@
Similarly, by comparing the scenarios generated with different levels of risk sensitivity (from left to right in Figure 7), we find that imposing a higher confidence level leads to more conservative and cautious driving behaviors. A higher confidence level forces the agent to satisfy the constraints with greater probability, resulting in driving strategies that feature lower speeds, increased spacing between vehicles, and more careful navigation through complex traffic scenarios. On the other hand, setting a lower confidence level results in more aggressive and risk-seeking driving behaviors, characterized by faster vehicle speeds and shorter following distances. The system’s tolerance for some aggressive behaviors (such as overtaking etc.) also increases. This approach can lead to a higher risk of collisions and increase the likelihood of critical safety scenarios in traffic.
@@
These diverse scenarios are crucial for investigating the trade-offs between aggressive and conservative driving behaviors. By analyzing AV policies within these scenarios, we can more effectively evaluate the optimality of driving strategies across different levels of competition and congestion. Furthermore, the comprehensive 2D and 3D simulations provided by our TrafficGamer  from both the third-person and first-person perspectives offer a detailed understanding of traffic dynamics. These simulations demonstrate how variations in risk sensitivity and distance constraints impact vehicle behavior in real-world driving scenarios.
@@
Figure 7: Visualization of generated trajectories with high diversity (Argoverse2). Visualization of all Argoverse2 scenarios in a 3x3 grid layout with 2D and 3D simulations. In each grid, the distance constraint relaxes from top to bottom, and the risk level strengthens from left to right. In the 2D scenes, yellow cars are algorithm-controlled, green cars and gray cars are environmental vehicles, black lines are vehicle trajectories, and blue arrows indicate travel direction. In the 3D scenes, red cars are algorithm-controlled, while blue cars represent environmental vehicles.
@@
Figure 8: Visualization of generated trajectories with high diversity (Waymo). Visualization of all Waymo scenarios in a 3x3 grid layout with 2D and 3D simulations. In each grid, the distance constraint relaxes from top to bottom, and the risk level strengthens from left to right. In the 2D scenes, yellow cars are algorithm-controlled, green cars and gray cars are environmental vehicles, black lines are vehicle trajectories, and blue arrows indicate travel direction. In the 3D scenes, red cars are algorithm-controlled, while other cars represent environmental vehicles.
Figure 9: Variation in Lagrangian parameters under varying constraints. Each row represents a scenario and scenarios 1-6 correspond to Y-Junction, Dense-lane intersection, Roundabout, Dual-lane intersection, T-Junction, and Merge respectively. Each column corresponds to one of the agents in the multi-agent environment. Because of the space limitations, the Lagrangian parameter results of other agents can be seen in Supplementary Section 3b.
@@ Discussion
We have demonstrated that our model, TrafficGamer, can effectively represent various safety-critical traffic scenarios by capturing a range of CCEs. To the best of our knowledge, this is the first algorithm able to fine-tune generative world models to accurately model competitive and collaborative behaviors among multiple agents across various degrees of constraints and dynamic environments. Most importantly, TrafficGamer can accurately characterize traffic congestion scenarios that are frequently observed in reality but are underrepresented in datasets, such as roundabouts, intersections, and merge points. This capability ensures the high fidelity of TrafficGamer in simulating real-world applications.
To facilitate the reliable generation of our safety-critical scenarios, we resolved three critical challenges, including (1) how to guarantee the fidelity of generated trajectories, (2) how to efficiently capture the CCE of each scenario by modeling the competition behaviors of vehicles, and (3) how to dynamically adapt the strength of the equilibrium. These safety-critical scenarios can serve as important testbeds for AV when evaluating the reliability and robustness of AV-control algorithms before their practical deployment. Traditional simulation systems have relied heavily on manually designed rules or data-driven trajectory imitation, often resulting in scenarios that lack fidelity and diversity. Our model, TrafficGamer, addresses these limitations by generating a variety of realistic, safety-critical scenarios. It is important to note that our method currently models vehicle behaviors on a static map. Future enhancements for TrafficGamer may include the integration of multi-modal large models to generate scenarios that incorporate variations in vehicle behaviors in response to environmental factors, such as weather conditions, the time of day, and topography.